# Graph neural network outputs are almost surely asymptotically constant

### v1

@@ Appendix C Sparse Erdős-Rényi graphs
    For a $(k+1)$-tuple $(\bar{u},v)$ define:
    $f_{\pi}^{\circ}(v;\bar{u})\coloneqq e_{\rho}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|}))h(e_{\eta}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|})))$
    where $t^{\prime}=\mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)$ and $s_{1},\ldots,s_{|t^{\prime}|}$ is the canonical enumeration of $t^{\prime}$.
    Similarly define:
    $g_{\pi}^{\circ}(v;\bar{u})\coloneqq h(e_{\eta}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|})))$
    Since $f_{\pi}^{\circ}$ is bounded, for every $\gamma>0$, by Hoeffding’s Inequality (Theorem A.4), for any tuple $\bar{u}$ and $t^{\prime}\in S$ we have that:
    $\mathbb{P}\left(\left\lVert\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}f_{\pi}^{\circ}(v;\bar{u})-|\llbracket t^{\prime}(\bar{u})\rrbracket|\cdot\operatorname*{\mathbb{E}}_{v}\left[f_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]\right\rVert\geq|\llbracket t^{\prime}(\bar{u})\rrbracket|\gamma\right)$
    $\leq 2\exp\left(-\frac{2\llbracket t^{\prime}(\bar{u})\rrbracket\gamma^{2}}{(b-a)^{2}}\right)$
    $\leq 2\exp\left(-\frac{2n(q_{t}^{\prime}-\epsilon)\gamma^{2}}{(b-a)^{2}}\right)$
@@
    $\left\lVert\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}f_{\pi}^{\circ}(v;\bar{u})-|\llbracket t^{\prime}(\bar{u})\rrbracket|\cdot\operatorname*{\mathbb{E}}_{v}\left[f_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]\right\rVert<|\llbracket t^{\prime}(\bar{u})\rrbracket|\gamma$
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}f_{\pi}^{\circ}(v;\bar{u})-q_{t^{\prime}}\operatorname*{\mathbb{E}}_{v}\left[f_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]\right\rVert<\epsilon$ ($\triangle_{f}$)
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}g_{\pi}^{\circ}(v;\bar{u})-q_{t^{\prime}}\operatorname*{\mathbb{E}}_{v}\left[g_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]\right\rVert<\epsilon$ ($\triangle_{g}$)
    By the induction hypothesis, there is $N_{6}$ such that for all $n\geq N_{6}$ with probability at least $1-\theta$, for every $t^{\prime}\in S$ and every $(k+1)$-tuple of nodes $(\bar{u},v)$ such that $\mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}$ we have that both:
    $\left\lVert e_{\rho}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|}))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\rho}$)
@@
    $\left\lVert e_{\eta}^{t^{\prime}}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t^{\prime}|}))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\eta}$)
@@
    $\left\lVert\frac{\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)}{\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)}-\frac{f_{\pi}^{t}(\mathrm{H}(u_{1}),\ldots,\mathrm{H}(u_{k}))}{g_{\pi}^{t}(\mathrm{H}(u_{1}),\ldots,\mathrm{H}(u_{k}))}\right\rVert<\iota$
@@
    $\left\lVert\frac{1}{n}\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)-f_{\pi}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))\right\rVert<\zeta$ (11)
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}(\mathrm{H}(s_{1}),\ldots,\mathrm{H}(s_{|t|}))\right\rVert<\xi$ (12)
@@
    $\left\lVert h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}^{\circ}(v;\bar{u})\right\rVert<\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(v;\bar{u})\right\rVert<\epsilon$
@@
    $\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(v;\bar{u})=\frac{1}{n}\sum_{t^{\prime}\in\mathsf{IsoTp}_{k,\ell^{\prime}}}\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}g_{\pi}^{\circ}(v;\bar{u})$
@@
    $\left\lVert\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(v;\bar{u})-\frac{1}{n}\sum_{t^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}g_{\pi}^{\circ}(v;\bar{u})\right\rVert<\epsilon\Lambda$
@@
    $\left\lVert\frac{1}{n}\sum_{t^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \mathrm{IsoTp}_{\ell^{\prime}}(\bar{u},v)=t^{\prime}\end{subarray}}g_{\pi}^{\circ}(v;\bar{u})-\frac{1}{n}\sum_{t^{\prime}\in S}q_{t^{\prime}}\operatorname*{\mathbb{E}}_{v}\left[g_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]\right\rVert\leq\epsilon$
    Now note that by the independence of the node features:
    $\operatorname*{\mathbb{E}}_{v}\left[g_{\pi}^{\circ}(v;\bar{u})\mid t^{\prime}(\bar{u},v)\right]=\operatorname*{\mathbb{E}}_{\bar{\bm{b}}}\left[h(e_{\eta}^{t^{\prime}}({\bm{a}}_{1},\ldots,{\bm{a}}_{k},{\bm{b}}_{k+1},\ldots,{\bm{b}}_{|t^{\prime}|}))\right]$
@@
    $\left\lVert\frac{1}{n}\sum_{t^{\prime}\in S}q_{t^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}}\left[h(e_{\eta}^{t^{\prime}}({\bm{a}}_{1},\ldots,{\bm{a}}_{k},{\bm{b}}_{k+1},\ldots,{\bm{b}}_{|t^{\prime}|}))\right]-g_{\pi}^{t}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\right\rVert<\epsilon\Lambda$
@@
Application to prove the final theorem, Theorem 5.1 for the sparse case.
To complete the proof, we note that there is only one $(0,\ell)$ local isomorphism type, namely $t=(\varnothing,\varnothing)$.
We have that $e_{\tau}^{t}$ is a constant $z$, and hence by the induction hypothesis applied to it, for every $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
$\left\lVert\llbracket\tau\rrbracket-z\right\rVert<\epsilon$
Appendix D Proof for the stochastic block model
We now prove the convergence result for the stochastic block model, Theorem 5.2. The proof follows the same structure as the density case in Theorem 5.1, except that the notion of graph type is augmented slightly. We therefore only indicate the differences in the proof.
@@ Appendix D Proof for the stochastic block model
To prove the result, we first need that the random work positional encodings converge, as in Lemma B.5. In fact they converge to $\bm{0}$.
Lemma D.1.
Let $n_{1},\ldots,n_{m}\colon\mathbb{N}\to\mathbb{N}$ and ${\bm{P}}$ be as in LABEL:thm:SBM_convergenceAgg[sum*]. Then for every $k\in\mathbb{N}$ and $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for all nodes $v$ we have that:
[…]
@@
Note that $nr_{j}$ is the expected degree of a node in community $j$. By Hoeffding’s Inequality (Theorem A.4) and a union bound, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$, for all $v$ we have that:
@@
So assume $r_{C(v)}>0$. Let $j=C(v)$. Then as in the proof of Lemma B.5 we can show that the proportion of random walks of length $k$ starting at $v$ is at most:
@@
We then proceed as in the proof of Theorem 5.1. The only difference is that when showing equation ($\triangle_{f}$) and equation ($\triangle_{g}$) we use the fact that the expected proportion of type extensions $t^{\prime}(\bar{u},v)$ of a type $t(\bar{u})$ at a node $u_{j}$ is $\alpha(t,t^{\prime})$.

### v2

@@ D.1 Combinatorial tools for the sparse case
    Define the function:
    $f_{\pi}^{\circ}(\bar{\bm{a}},\bar{\bm{b}})\coloneqq e_{\rho}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}}))\cdot h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))$
    and similarly:
    $g_{\pi}^{\circ}(\bar{\bm{a}},\bar{\bm{b}})\coloneqq h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))$
    Note that these are just the integrands used in $f_{\pi}^{T}(\bar{\bm{a}})$ and $g_{\pi}^{T}(\bar{\bm{a}})$.
    Now, for any $(k+1)$-tuple of nodes $(\bar{u},v)$ let $\bar{\bm{a}}^{\bar{u}}$ be the tuple of node features of $\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}-1}(\bar{u})$ and $\bar{\bm{b}}^{(\bar{u},v)}$ be the tuple of node features of the remaining nodes in $\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)$, ordered as in the canonical enumeration of $\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)$.
    Note that $f_{\pi}^{\circ}$ is bounded, say by $\Lambda$. Then for every $\gamma>0$, by Hoeffding’s Inequality (Theorem B.4) for any tuple $\bar{u}$ and $T^{\prime}\in S$ we have that:
    $\mathbb{P}\left(\left\lVert\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-|\llbracket T^{\prime}(\bar{u})\rrbracket|\cdot\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert\geq|\llbracket T^{\prime}(\bar{u})\rrbracket|\gamma\right)$
    $\leq 2\exp\left(-\frac{2\llbracket T^{\prime}(\bar{u})\rrbracket\gamma^{2}}{\Lambda^{2}}\right)$
    $\leq 2\exp\left(-\frac{2n(q_{T}^{\prime}-\epsilon)\gamma^{2}}{\Lambda^{2}}\right)$
@@
    $\left\lVert\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-|\llbracket T^{\prime}(\bar{u})\rrbracket|\cdot\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<|\llbracket T^{\prime}(\bar{u})\rrbracket|\gamma$
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{f}$)
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{g}$)
    By the induction hypothesis, there is $N_{6}$ such that for all $n\geq N_{6}$ with probability at least $1-\theta$, for every $T^{\prime}\in S$ and every $(k+1)$-tuple of nodes $(\bar{u},v)$ such that $\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}$ we have that both:
    $\left\lVert e_{\rho}^{T^{\prime}}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T^{\prime}|}))-\llbracket\rho(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\rho}$)
@@
    $\left\lVert e_{\eta}^{T^{\prime}}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T^{\prime}|}))-\llbracket\eta(\bar{u},v)\rrbracket\right\rVert<\epsilon$ ($\dagger_{\eta}$)
@@
    $\left\lVert\frac{\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)}{\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)}-\frac{f_{\pi}^{T}(H_{G}(u_{1}),\ldots,H_{G}(u_{k}))}{g_{\pi}^{T}(H_{G}(u_{1}),\ldots,H_{G}(u_{k}))}\right\rVert<\iota$
@@
    $\left\lVert\frac{1}{n}\sum_{v}\llbracket\rho(\bar{u},v)\rrbracket h(\llbracket\eta(\bar{u},v)\rrbracket)-f_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))\right\rVert<\zeta$ (6)
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}^{T}(H_{G}(s_{1}),\ldots,H_{G}(s_{|T|}))\right\rVert<\xi$ (7)
@@
    $\left\lVert h(\llbracket\eta(\bar{u},v)\rrbracket)-g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})\right\rVert<\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{v}h(\llbracket\eta(\bar{u},v)\rrbracket)-\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})\right\rVert<\epsilon$
@@
    $\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})=\frac{1}{n}\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})$
@@
    $\left\lVert\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-\frac{1}{n}\sum_{T^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})\right\rVert<\epsilon\Lambda$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-\frac{1}{n}\sum_{T^{\prime}\in S}q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert\leq\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]-g_{\pi}^{T}(\bar{\bm{a}}^{\bar{u}})\right\rVert<\epsilon\Lambda$
@@
Application to prove the final theorem, Theorem 5.1 for the sparse Erdős-Rényi and Barabási-Albert cases.
To complete the proof, we note that the term $\tau$ has no free variables, and as a subterm of itself has reach 00. The controller $e_{\tau}^{\varnothing}$ of is a constant ${\bm{z}}$, and hence by the induction hypothesis applied to it, for every $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$ we have that:
$\left\lVert\llbracket\tau\rrbracket-{\bm{z}}\right\rVert<\epsilon$
Appendix E Proof for the stochastic block model
We now prove the convergence result for the stochastic block model. The proof follows the same structure as the density case in Theorem 5.1, except that the notion of graph type is augmented slightly. We therefore only indicate the differences in the proof.
@@ Appendix E Proof for the stochastic block model
To prove the result, we first need that the random work positional encodings converge, as in Lemma C.6. In fact they converge to $\bm{0}$.
Lemma E.1.
Let $n_{1},\ldots,n_{m}\colon\mathbb{N}\to\mathbb{N}$ and ${\bm{P}}$ be as in Theorem 5.1. Then for every $k\in\mathbb{N}$ and $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for all nodes $v$ we have that:
$\left\lVert\text{rw}_{k}(v)\right\rVert<\epsilon$
@@
Note that $nr_{j}$ is the expected degree of a node in community $j$. By Hoeffding’s Inequality (Theorem B.4) and a union bound, there is $N$ such that for all $n\geq N$ with probability at least $1-\theta$, for all $v$ we have that:
@@
So assume $r_{C(v)}>0$. Let $j=C(v)$. Then as in the proof of Lemma C.6 we can show that the proportion of random walks of length $k$ starting at $v$ is at most:
@@
$\left\lVert e_{\pi}(H_{G}(\bar{u}),\mathrm{GrTp}(\bar{u}))-\llbracket\pi(\bar{u})\rrbracket\right\rVert<\epsilon$
We then proceed as in the proof of the non-sparse Erdős-Rényi cases of Theorem 5.1. The only difference is that when showing equation ($\triangle_{f}$) and equation ($\triangle_{g}$) we use the fact that the expected proportion of type extensions $t^{\prime}(\bar{u},v)$ of a type $t(\bar{u})$ at a node $u_{j}$ is $\alpha(t,t^{\prime})$.
Appendix F Additional experiments
In this section we provide additional experiments using MeanGNN, GCN [30], GAT [52], and GPS+RW with random walks of length up to 5, over the distributions $\mathrm{ER}(n,p(n)=0.1)$, $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$, $\mathrm{ER}(n,p(n)=\frac{1}{50n})$ and $\mathrm{BA}(n,m=5)$. We also experiment with an SBM of 10 communities with equal size, where an edge between nodes within the same community is included with probability $0.7$ and an edge between nodes of different communities is included with probability $0.1$.
Setup. Our setup is carefully designed to eliminate confounding factors:
  1. We consider 5 models with the same architecture, each having randomly initialized weights, utilizing a $\operatorname{ReLU}$ non-linearity, and applying a softmax function to their outputs. Each model uses a hidden dimension of $128$, $3$ layers and an output dimension of $5$.
  2. We draw graphs of sizes up to 10,000, where we take 100 samples of each graph size. Node features are independently drawn from $U[0,1]$ and the initial feature dimension is $128$.
Much like in Section 6, the convergence of class probabilities is apparent across all models and graph distributions, in accordance with our main theorems (Q1). We again observe that attention-based models such as GAT and GPS+RW exhibit delayed convergence and greater standard deviation in comparison to MeanGNN and GCN, which further strengthening our previous conclusions.
Figure 5: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=0.1)$ graph distributions, as we draw increasingly larger graphs.
Figure 6: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ graph distributions, as we draw increasingly larger graphs.
Figure 7: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{1}{50n})$ graph distributions, as we draw increasingly larger graphs.
Figure 8: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the SBM graph distributions, as we draw increasingly larger graphs.
Figure 9: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{BA}(n,m=5)$ graph distributions, as we draw increasingly larger graphs.